### v2

Abstract
Zero-shot anomaly detection (ZSAD) enables identifying and localizing defects in unseen categories by relying solely on generalizable features rather than requiring any labeled examples of anomalies. However, existing ZSAD methods, whether using fixed or learned prompts, struggle under domain shifts because their training data are derived from limited training domains and fail to generalize to new distributions. In this paper, we introduce PILOT, a framework designed to overcome these challenges through two key innovations: (1) a novel dual-branch prompt learning mechanism that dynamically integrates a pool of learnable prompts with structured semantic attributes, enabling the model to adaptively weight the most relevant anomaly cues for each input image; and (2) a label-free test-time adaptation strategy that updates the learnable prompt parameters using high-confidence pseudo-labels from unlabeled test data. Extensive experiments on 13 industrial and medical benchmarks demonstrate that PILOT achieves state-of-the-art performance in both anomaly detection and localization under domain shift.
Accepted at BMVC 2025.
1 Introduction
Anomaly detection and localization is critical in safety-critical applications, including industrial inspection and medical diagnostics, where timely identification of rare deviations from normal patterns can prevent costly failures and improve patient outcomes. Traditional anomaly detection methods include supervised approaches that require labeled anomalous examples and unsupervised techniques relying solely on normal data [Bozic et al.(2021)Bozic, Tabernik, and Skocaj, Hojjati and Armanfard(2023), Li et al.(2021)Li, Sohn, Yoon, and Pfister, Roth et al.(2022)Roth, Pemula, Zepeda, Schölkopf, Brox, and Gehler, Karami et al.(2025)Karami, Ho, and Armanfard, Lai et al.(2023)Lai, Ho, and Armanfard]. Supervised methods require scarce and diverse anomaly labels, whereas unsupervised approaches avoid any need for anomalous samples by modeling only normal data, though they depend on comprehensive, representative normal data from the target domain [Chandola et al.(2009)Chandola, Banerjee, and Kumar]. Zero-shot Anomaly Detection (ZSAD) has emerged as a powerful approach by leveraging pretrained vision–language models (VLMs) such as CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.], which learns joint representations of images and text from large-scale image–text data, to identify anomalies in visual data without requiring any target-domain annotations of normal or anomalous samples.
Early successful ZSAD approaches utilized CLIP by designing explicit textual prompts that described both the object category and its condition, such as “normal bottle” or “anomalous transistor”. By crafting a large number of prompts that described relevant combinations of object category and condition, these methods enabled zero-shot classification and significantly outperformed vanilla CLIP baselines [Jeong et al.(2023)Jeong, Zou, Kim, Zhang, Ravichandran, and Dabeer, Chen et al.(2024)Chen, Zhang, Tian, He, Zhang, Wang, Wang, and Liu, Esmaeilpour et al.(2022)Esmaeilpour, Liu, Robertson, and Shu]. These approaches are known as fixed prompt methods because they rely on manually designed textual prompts. More recent methods have shifted towards a single learnable prompt, where the prompt itself is represented as a trainable embedding and optimized during training. This optimization is performed using auxiliary data, which refers to anomaly detection datasets that are separate from the target domain. For example, AnomalyCLIP [Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen] learns single object agnostic prompt embeddings to capture generic anomaly patterns, while AdaCLIP [Cao et al.(2024)Cao, Zhang, Frittoli, Cheng, Shen, and Boracchi] further integrates visual knowledge into the prompt embeddings to enhance model adaptation.
Despite these advances, both fixed and single learnable prompt approaches remain vulnerable to domain shift, which refers to performance degradation when data distribution in the target domain differs significantly from training data. Although pretrained VLMs like CLIP have strong generalization ability, their performance can drop sharply when there are significant differences between the data seen during pretraining and the target environment [Bose et al.(2024)Bose, Jha, Fini, Singha, Ricci, and Banerjee, Feng et al.(2024)Feng, Yu, Jin, Yu, Xiao, and Chen]. Even when learnable prompt embeddings are further trained on auxiliary anomaly detection datasets, that are different from the actual deployment domain, these models often struggle to represent the full diversity of possible anomalies or adapt to new categories encountered at test time. This limited generalization occurs for two main reasons: first, the auxiliary and target datasets may have significant distribution differences, as we empirically demonstrate in Appendix B.2; second, single prompt embeddings risk overfitting to the patterns and features specific to the auxiliary dataset, which further reduces their ability to handle unseen anomalies. As a result, prompts derived from narrowly defined auxiliary sources may fail to capture diverse visual features in the target data [Marzullo and Ranzini(2024)]. Therefore, to address these challenges, we introduce a framework that learns and dynamically weights multiple prompts, rather than relying on a fixed or single learnable prompt.
On the other hand, Test-Time Adaptation (TTA) techniques have demonstrated improved model robustness in various computer vision tasks by refining model parameters on unlabeled test data [Nado et al.(2020)Nado, Padhy, Sculley, D’Amour, Lakshminarayanan, and Snoek, Liang et al.(2020)Liang, Hu, and Feng, Schneider et al.(2020)Schneider, Rusak, Eck, Bringmann, Brendel, and Bethge]. Although TTA is designed to address domain shift between training and test data, applying naive TTA strategies – without tailoring them to the specific task at hand – can inadvertently degrade performance. When used blindly, especially in tasks such as detection or localization, the adaptation process may optimize surrogate objectives that are misaligned with the true goals, often leading to suboptimal outcomes [Sun et al.(2020)Sun, Wang, Liu, Miller, Efros, and Hardt, Liu et al.(2021)Liu, Kothari, Van Delft, Bellot-Gurlet, Mordan, and Alahi, Liang et al.(2025)Liang, He, and Tan]. AnovL [Deng et al.(2023)Deng, Zhang, Bao, and Li] was among the first to introduce TTA for anomaly localization refinement in ZSAD scenarios, but its focus is limited to localization and relies on fixed prompts. These limitations motivate us to explore how both anomaly detection and localization can be improved under domain shift by effectively utilizing unlabeled target data.  [^1] We present the Related Works section in Section A of the Supplementary Material..
To address these challenges, we propose a novel framework, Prompt Learning with Integrated Label-Free Test-Time Adaptation for ZerO-ShoT Anomaly Detection (PILOT), a framework designed to mitigate the limitations of current ZSAD methods under domain shift. Instead of relying on a single learnable prompt, PILOT constructs a pool of learnable prompts, allowing the prompts to adaptively learn from various anomalies present in the auxiliary dataset. This design enables the model to adaptively weight the most relevant prompts for each input image, allowing for more nuanced representations of diverse anomalies and supporting more effective TTA. During inference, our label-free test-time adaptation strategy updates all learnable prompts according to their alignment with each test image, using high-confidence pseudo-labels generated from unlabeled target data. This targeted adaptation facilitates rapid and reliable adjustment to new domains (i.e., the test domain) without requiring any labeled data. In addition, to mitigate overfitting to the auxiliary dataset and enhance adaptation stability, PILOT incorporates an attribute memory bank, which consists of a curated set of fixed prompts that describe normal and anomalous states, leveraging pretrained semantic knowledge from the text encoder. This attribute memory bank serves as a semantic anchor, helping to stabilize localization performance during TTA. Together, these innovations enable PILOT to achieve robust anomaly detection and localization across a variety of challenging real-world scenarios. Our contributions are summarized as follows:
  • We propose a dual-branch prompt learning framework that dynamically weight the most relevant prompts from both learnable prompt pool and attribute memory bank, enabling the model to leverage the most relevant anomaly cues for each input image.
  • To our knowledge, this is the first label-free TTA strategy in ZSAD that leverages high-confidence pseudo-labels from unlabeled test samples to rapidly adapt prompt parameters at inference, yielding improvements in both anomaly detection and localization under domain shift.
  • Extensive evaluation on industrial and medical benchmarks shows that our method outperforms existing approaches in both anomaly detection and localization metrics.
2 Proposed Method
ZSAD involves detecting and localizing anomalies in images from categories not encountered during the training phase. Formally, in the training phase, we employ an auxiliary dataset defined as $\mathcal{D}_{\mathrm{aux}}={(I_{i},y_{i},G_{i})}_{i=1}^{|\mathcal{D}_{\mathrm{aux}}|}$, where each image $I_{i}\!\in\!\mathbb{R}^{H\times W\times 3}$ is paired with an image-level label $y_{i}\!\in\!\{0,1\}$ (0 denotes normal and 1 denotes anomalous) and a pixel-level anomaly mask
$G_{i}\!\in\!\{0,1\}^{H\times W}$. In the inference phase, we consider an unlabeled target dataset $\mathcal{D}_{\mathrm{tar}}=\bigl{\{}I_{t}\bigr{\}}_{t=1}^{|\mathcal{D}_{\mathrm{tar}}|},$ where each $I_{t}\in\mathbb{R}^{H\times W\times 3}$. For every $I_{t}$, the model outputs an anomaly score $\hat{y}_{t}\in[0,1],$ (with values closer to 1 indicating higher abnormality likelihood), along with a $\hat{G}_{t}\in[0,1]^{H\times W}$ that spatially localizes potential anomalous regions at pixel level. Crucially, the category sets of $\mathcal{D}_{\mathrm{aux}}$ and $\mathcal{D}_{\mathrm{tar}}$ are strictly disjoint, ensuring that the target domain remains unseen during training and thus establishing a true zero-shot setting.
To address the limitations of existing approaches, we propose a dual-branch framework that combines a learnable prompt pool $\mathcal{P}$ with an attribute memory bank $\mathcal{U}$. The training phase of our PILOT framework, depicted in Fig. 1, highlights these two key components. Further details on the inference phase are provided in Section 2.5, with in-depth descriptions of each component in the following sections and the complete pseudocode in Appendix E.
Figure 1: Overview of PILOT’s training phase. The left panel shows the main workflow. The right panel details the Learnable Prompt Pool $\mathcal{P}$ and Attribute Memory Bank $\mathcal{U}$.
2.1 Learnable Prompt Pool
A successful recent approach for CLIP-based ZSAD employs a single learnable prompt, defined as a pair of sequences $(p^{n},p^{a})$ representing normal and anomalous states, respectively, for binary anomaly classification [Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen]. Formally, this prompt can be represented as:
$p^{n}=[v^{1},\!\dots,\!v^{L},\!\text{``[normal state] object''}],\;p^{a}=[v^{1},\!\dots,\!v^{L},\!\text{``[anomalous state] object''}]$ (1)
where $L$ is the number of learnable prefix vectors, and each $v^{(i)}\in\mathbb{R}^{d}$ is a $d$-dimensional vector trained jointly with the model. These vectors form a “soft prompt,” intuitively acting as special tokens whose values steer the model’s attention without modifying its core parameters, follows the prefix-tuning paradigm, which guides VLMs while keeping the main model weights frozen [Li and Liang(2021), Khattak et al.(2023)Khattak, Rasheed, Maaz, Khan, and Khan]. The placeholders “[normal state]” and “[anomalous state]” filled with descriptors of the object’s condition, such as “flawless” and “damaged.”
Despite its simplicity and effectiveness, relying solely on a single learnable prompt for all input images limits generalization under domain shift, as discussed in the introduction, because a single pair of prompts often fails to adequately represent the diversity of anomaly patterns encountered in different scenarios. To address this limitation, we propose extending this single prompt to a learnable prompt pool, containing multiple learnable prompts, each associated with additional embeddings for adaptive weighting and alignment. Formally, the proposed prompt pool is defined as $\mathcal{P}=\{(p_{k}^{n},\,p_{k}^{a},\,s_{k},\,r_{k})\}_{k=1}^{K}$, where $K=|\mathcal{P}|$ is the cardinality of the prompt pool. Each $(p_{k}^{n},p_{k}^{a})$ is constructed similarly to previous works, and the additional learnable embeddings $s_{k}$ and $r_{k}$ enable adaptive weighting of the prompts and alignment with the input image features. Here, $s_{k}$ is a modulation vector for prompt $k$, applied element-wise to the image feature $x_{\mathrm{cls}}$ to adaptively emphasize information relevant to that prompt, while $r_{k}$ provides a prompt-specific reference for measuring similarity. These embeddings facilitate an adaptive emphasis on prompts that best align with the diverse patterns present in the auxiliary dataset. To achieve this adaptive prompt weighting, we introduce a prompt-querying function, which computes prompt relevance scores given an image feature $x_{\mathrm{cls}}\in\mathbb{R}^{d}$ produced by the CLIP vision encoder, i.e. $f_{\mathrm{visual}}$, from an image $I_{i}$ in $\mathcal{D}_{\mathrm{aux}}$,:
$Q_{\mathrm{lrn}}\bigl{(}x_{\mathrm{cls}},\,\mathcal{P}\bigr{)}=\bigl{\{}\alpha_{k}\coloneqq\mathrm{sim}\bigl{(}x_{\mathrm{cls}}\odot s_{k},\,r_{k}\bigr{)}\bigr{\}}_{k=1}^{K},$ (2)
where $\odot$” denotes element-wise multiplication, and $\mathrm{sim}(\cdot,\cdot)$ represents cosine similarity. Intuitively, $\alpha_{k}$ quantifies how well the image content aligns with the k-th prompt pair; this is realized by weighting the image feature $x_{\mathrm{cls}}$ with the attention mask $s_{k}$ and measuring its similarity to the associated reference embedding $r_{k}$. The higher the score $\alpha_{k}$ is, the more similar the input image is to the k-th prompt $(p_{k}^{n},\,p_{k}^{a})$. These scores are then used to form weighted combinations of the prompt token sequences as $\bar{p}^{n}=\sum_{k=1}^{K}\alpha_{k}p_{k}^{n}$, $\bar{p}^{a}=\sum_{k=1}^{K}\alpha_{k}p_{k}^{a}$. The resulting $\bar{p}^{n}$ and $\bar{p}^{a}$ are subsequently encoded using the CLIP pretrained text encoder to obtain the final text embeddings $T_{n}=f_{\text{text}}(\bar{p}^{n})$ and $T_{a}=f_{\text{text}}(\bar{p}^{a})$, which are employed for downstream anomaly scoring and localization. This process is visualized in Figure 1.
Although the text encoder $f_{\mathrm{text}}$ is frozen to preserve CLIP’s rich pretrained textual representations, we employ a prefix-tuning strategy to enable effective fine-tuning and adaptation of CLIP. Inspired by similar prefix-tuning methods [Khattak et al.(2023)Khattak, Rasheed, Maaz, Khan, and Khan], we introduce additional learnable parameters to $f_{\mathrm{text}}$. More specifically, as is shown in Figure 1, layer-wise prefix vectors are defined as $\bigl{\{}\,w_{j}^{(\ell)}\in\mathbb{R}^{d}\mid j=1,\dots,J,\;\ell=1,\dots,L^{\prime}\bigr{\}}$, where $f_{\mathrm{text}}$ consists of a total of $J^{\prime}$ transformer layers, and $J\leq J^{\prime}$ denotes the number of initial layers to which learnable vectors are appended. $L^{\prime}$ represents the number of learnable vectors per layer. Specifically, these prefix vectors are prepended as additional tokens to the input sequence at the first $J$ transformer layers, leaving the original transformer layers unchanged. At each layer $j$, these vectors $\{w_{j}^{(\ell)}\}_{\ell=1}^{L^{\prime}}$ (illustrated as green blocks within the $f_{\mathrm{text}}$ module of Figure 1) are added before the standard, frozen token embeddings (represented as gray blocks in the the $f_{\mathrm{text}}$ module of Figure 1) prior to the self-attention mechanism in in each frozen transformer layer. This approach enables adaptation of the model’s internal representations without modifying the pretrained CLIP parameters, thereby preserving their robustness and generalization capabilities. Furthermore, as is illustrated in Figure 1, PILOT fine-tunes only the final $d$-dimensional projection layer within the vision encoder $f_{\mathrm{visual}}$. By restricting updates exclusively to this projection layer, we substantially reduce memory overhead and computational complexity. This targeted fine-tuning strategy also mitigates the risk of overfitting to auxiliary datasets, maintaining the strong generalization properties inherent to CLIP [Fahes et al.(2024)Fahes, Vu, Bursuc, Pérez, and de Charette].
2.2 Attribute Memory Bank
Although the learnable prompt pool $\mathcal{P}$ aims to provide adaptive and diverse representations tailored to the auxiliary dataset, it may still be susceptible to overfitting when deployed on target domains. To mitigate this issue, we further introduce a fixed attribute memory bank comprising attribute embeddings extracted from the pretrained CLIP text encoder $f_{\text{text}}$ (prior to applying our modifications). Recent works, such as WinCLIP [Jeong et al.(2023)Jeong, Zou, Kim, Zhang, Ravichandran, and Dabeer], have demonstrated the robustness of fixed textual prompts by encoding multiple natural-language templates. Inspired by this, we construct our attribute memory bank from a set of descriptive natural-language templates (e.g., “A bright photo of a [State] object”), where [State] indicates descriptive attributes of the object’s condition (e.g., “normal", “corroded", or “contaminated").
Each attribute template is passed through the frozen CLIP text encoder $f_{\mathrm{text}}$, without any additional learnable tokens $w_{j}^{(\ell)}$, to produce a fixed attribute embedding, formally defined as $\mathcal{U}=\bigl{\{}(u_{c},\,g_{c})\bigr{\}}_{c=1}^{C},$ where $C$ is the number of attribute embeddings, and each $u_{c}\in\mathbb{R}^{d}$ represents the EOS-token embedding of the corresponding template. Unlike WinCLIP, which averages embeddings across multiple prompts, we associate each attribute embedding $u_{c}$ with a learnable modulation vector $g_{c}$, similar to the learnable prompt pool. Similar to how $s_{k}$ adapts the image features for each prompt, $g_{c}$ dynamically modulates the contribution of each attribute according to the input image. Specifically, given an image feature $x_{\mathrm{cls}}$, we introduce the attribute-querying function to compute attribute relevance scores as below:
$Q_{\mathrm{att}}\bigl{(}x_{\mathrm{cls}},\,\mathcal{U}\bigr{)}=\bigl{\{}\beta_{c}\coloneqq\mathrm{sim}\bigl{(}x_{\mathrm{cls}}\odot g_{c},\,u_{c}\bigr{)}\bigr{\}}_{c=1}^{C},$ (3)
where $\odot$” denotes element-wise multiplication, and $\mathrm{sim}(\cdot,\cdot)$ is the cosine similarity. Here, $u_{c}$ itself serves as the reference embedding for each attribute, as it is obtained directly from the pretrained $f_{\mathrm{text}}$ and already contains rich semantic information. Each score $\beta_{c}$ quantifies the alignment between the image and the $c$-th attribute embedding. The resulting scores $\{\beta_{c}\}_{c=1}^{C}$ quantify the alignment between the input and each attribute. Following [Jang et al.(2016)Jang, Gu, and Poole], we inject i.i.d. Gumbel noise into the attribute logits and apply the $\operatorname{entmax}{1.5}$ operator [Peters et al.(2019)Peters, Niculae, and Martins] to obtain a sparse yet differentiable weighting over the most prominent cues:
$\{\phi_{c}\}_{c=1}^{C}=\operatorname{entmax}{1.5}(\beta_{c}+\text{noise}),\quad\sum_{c=1}^{C}\phi_{c}=1,\quad\phi_{c}\geq 0.$ (4)
We then compute the attribute text embedding as a weighted combination. $T_{att}=\sum_{c=1}^{C}\phi_{c}u_{c},{}\\ T_{att}\in\mathbb{R}^{d}$ The injected noise promotes exploration and prevents premature convergence to a single attribute, allowing the model to capture diverse and complementary semantic cues. By providing stable, pretrained semantic cues, the attribute memory bank acts as a semantic anchor during TTA discussed in Section 2.5. This mechanism helps to counteracts the potential overfitting and instability of learnable prompts, ensuring the model retains reliable semantic references even as it adapts to new domains.
2.3 Prompt Aggregation via Orthogonal Projection
Following [Yang et al.(2021)Yang, He, Fan, Shi, Xue, Li, Ding, and Huang], to robustly leverage both the stable semantics from the attribute embedding $T_{att}$ and the adaptive semantics from the learnable prompts ($T_{n}$ for normal and $T_{a}$ for anomalous), we aggregate these representations using an orthogonal projection operation instead of simple averaging or concatenation. For each prompt embedding $T_{(\cdot)}$ (where $(\cdot)\in\{n,a\}$), we decompose it into components parallel and orthogonal to $T_{att}$:
$T_{(\cdot)}^{\mathrm{fus}}=T_{\mathrm{att}}+(T_{(\cdot)}-\mathrm{Proj}_{T_{\mathrm{att}}}(T_{(\cdot)})),\quad\mathrm{Proj}_{\eta}(\psi)=\frac{\eta^{\top}\psi}{\|\eta\|^{2}}\,\eta$ (5)
This construction anchors the fused embedding in the robust attribute space, while enriching it with only those cues from the learnable prompt that are complementary to the attribute. Geometrically, $T_{(\cdot)}^{\mathrm{fus}}$ preserves all the core semantics provided by $T_{att}$, and selectively incorporates novel information from $T_{(\cdot)}$ that lies in the orthogonal complement. Further analysis can be found in Appendix E.8. This parameter-free aggregation prevents redundancy and enhances robustness to distributional shift, especially during TTA, as shown in Table 11.
2.4 Training Objectives
To achieve both accurate anomaly detection and precise localization, we jointly optimize image-level classification and pixel-level segmentation objectives. Joint training allows the two tasks to reinforce each other, as global context aids localization while fine-grained cues improve classification. All training is performed on the auxiliary dataset $\mathcal{D}_{aux}$. Given the aggregated text embedding ${T}_{(\cdot)}^{fus}$ and the global visual feature $x_{\mathrm{cls}}$, we compute the anomaly score and corresponding image-level loss:
$\hat{y}=\sigma\bigl{(}\mathrm{sim}(x_{\mathrm{cls}},\,T_{(\cdot)}^{\mathrm{fus}})\bigr{)},\quad\mathcal{L}_{\mathrm{image}}=\mathrm{BCE}(\hat{y},\,y)$ (6)
where $\sigma(\cdot)$ denotes the sigmoid function, $\mathrm{sim}(\cdot,\cdot)$ is cosine similarity, and $\mathrm{BCE}$ is the binary cross-entropy loss comparing the predicted score $\hat{y}$ to the ground-truth image label $y\in\{0,1\}$. For pixel-level localization, following [Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen], we extract patch features $\{\,X^{(m)}\in\mathbb{R}^{H\times W\times d}\}_{m\in\mathcal{M}}$ from a set $\mathcal{M}$ of selected layers in $f_{\mathrm{visual}}$. Each $X^{(m)}$ is formed by reshaping the patch token sequence into a spatial grid of size $H\times W$ with $d$-dimensional embeddings. For each $m\in\mathcal{M}$, we compute an anomaly map using the aggregated text embedding:
$\hat{G}^{(m)}=\sigma\bigl{(}\mathrm{sim}(X^{(m)},\,T_{(\cdot)}^{\mathrm{fus}})\bigr{)},\quad\mathcal{L}_{\mathrm{pixel}}=\frac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}\bigl{[}\mathrm{Focal}(\hat{G}^{(m)},\,G)+\mathrm{Dice}(\hat{G}^{(m)},\,G)\bigr{]}$ (7)
where $\hat{G}^{(m)}\in[0,1]^{H\times W}$ is the predicted anomaly probability map for layer $s$, $G$ is the ground-truth pixel-level mask, $\mathrm{Focal}(\cdot,\cdot)$ denotes the focal loss [Lin et al.(2017)Lin, Goyal, Girshick, He, and Dollár], and $\mathrm{Dice}(\cdot,\cdot)$ denotes the Dice loss [Sudre et al.(2017)Sudre, Li, Vercauteren, Ourselin, and Jorge Cardoso]. This combination balances pixel-wise discrimination and overlap accuracy. The final training objective combines both terms $\mathcal{L}_{\mathrm{PILOT}}=\mathcal{L}_{\mathrm{image}}+\mathcal{L}_{\mathrm{pixel}}.$
2.5 Test-Time Adaptation
A central challenge for robust ZSAD, as outlined in the Introduction, is achieving reliable adaptation to target domains that may differ significantly from the auxiliary dataset. While fine-tuning on $\mathcal{D}_{\mathrm{aux}}$ can calibrate model representations, substantial domain shifts often cause conventional approaches to overfit or fail to generalize, especially in the absence of labeled data from the target domain. To overcome these obstacles, we propose a TTA strategy that enables the model to adjust to new domains using only unlabeled target data, without compromising the integrity of previously learned representations.
The core idea of our TTA strategy is to identify a subset of highly reliable predictions (high-confidence pseudo-labels) from the target dataset and leverage these predictions to update model parameters. Specifically, for each image $I_{t}$ in the unlabeled target dataset $\mathcal{D}{\mathrm{tar}}$, we compute the anomaly score as $\tilde{y}t=\sigma\big{(}\mathrm{sim}(x{\mathrm{cls}}^{t},,T{(\cdot)}^{\mathrm{fus}})\big{)}$, where $x_{\mathrm{cls}}^{t}$ is the CLIP visual feature for $I_{t}$ and $T_{(\cdot)}^{\mathrm{fus}}$ denotes the fused prompt embedding selected and aggregated for $I_{t}$ using model parameters trained on $\mathcal{D}_{\mathrm{aux}}$. We then identify the most reliable samples by selecting the top and bottom $\rho$-fractions of scores:
$A=\mathrm{Top}_{\lceil\rho B\rceil}\{\tilde{y}_{t}\}_{t=1}^{B},\qquad N=\mathrm{Bot}_{\lceil\rho B\rceil}\{\tilde{y}_{t}\}_{t=1}^{B},$ (8)
where $B=|\mathcal{D}_{\mathrm{tar}}|$ and $0<\rho<0.5$. Pseudo-labels are then defined by $\hat{y}_{t}=1$ if $t\in A$, and $\hat{y}_{t}=0$ if $t\in N$. Adaptation is performed by optimizing an image-level binary cross-entropy loss on the selected samples, i.e., $\mathcal{L}_{\mathrm{image}}=\mathrm{BCE}(\hat{y}_{i},\tilde{y}_{i})$ for $i\in A\cup N$.
Critically, during TTA, we update only the parameters associated with the learnable prompt pool, including the learnable prompts, their corresponding embeddings $s_{k}$ and $r_{k}$. All other model parameters, such as those in the attribute memory bank and the core CLIP architecture, remain fixed. The updates for each learnable prompt are weighted by the attention scores defined previously (see Section 2.1), meaning prompts with higher attention scores for a given input contribute more significantly to parameter updates. We limit adaptation strictly to the image-level objective because, as shown in our experiments (Section 3.2), pseudo-labels derived at the image level exhibit considerably higher reliability compared to those at the pixel level. Pixel-level pseudo-labeling can be adversely affected by spatial ambiguity and noise exacerbated by domain shifts. By focusing adaptation solely on image-level labels, we reduce the risk of propagating noisy pseudo-labels, ensuring stable, unsupervised adaptation in the target domain. Our experimental results demonstrate that applying such a strategy improves both pixel-level and image-level detection on the target domain.
3 Experiments
Benchmarks. To demonstrate the effectiveness of our method, we conduct experiments on 13 widely used benchmarks spanning industrial inspection and medical imaging. Specifically, we evaluate on six industrial datasets: BTAD [Mishra et al.(2021)Mishra, Verk, Fornasier, Piciarelli, and Foresti], VisA [Zou et al.(2022)Zou, Jeong, Pemula, Zhang, and Dabeer], MVTec AD [Bergmann et al.(2019)Bergmann, Fauser, Sattlegger, and Steger], DAGM [Wieler and Hahn(2007)], SDD [Tabernik et al.(2020)Tabernik, Sela, Skvarc, and Skocaj], and DTD-Synthetic [Aota et al.(2023b)Aota, Tong, and Okatani] and seven medical datasets, subdivided into detection-only (HeadCT [Kitamura(2018)], BrainMRI [Kanade and Gumaste(2015)], Br35H [Hamada(2020)]) and localization (ISIC [Codella et al.(2018)Codella, Gutman, Celebi, Helba, Marchetti, Dusza, Kalloo, Liopyris, Mishra, Kittler, et al.], CVC ColonDB [Tajbakhsh et al.(2015)Tajbakhsh, Gurudu, and Liang], CVC ClinicDB [Bernal et al.(2015)Bernal, Sánchez, Fernández-Esparrach, Gil, Rodríguez, and Vilariño], Kvasir [Jha et al.(2020)Jha, Smedsrud, Riegler, Halvorsen, De Lange, Johansen, and Johansen]). The medical datasets for anomaly localization contain only anomalous samples and lack normal data; since our TTA strategy requires normal samples, results are reported without TTA (see Appendix H). To ensure strict zero-shot evaluation, auxiliary datasets exclude all target data. More specifically, for a fair comparison, we followed the related literature [Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen, Cao et al.(2024)Cao, Zhang, Frittoli, Cheng, Shen, and Boracchi], and used MVTec AD [Bergmann et al.(2019)Bergmann, Fauser, Sattlegger, and Steger] as the auxiliary dataset when reporting the performance on all other datasets, while when reporting the performance on MVTec AD, we trained the model on VisA [Zou et al.(2022)Zou, Jeong, Pemula, Zhang, and Dabeer] as the auxiliary dataset.
Evaluation Metrics. Following standard practices in zero-shot anomaly detection literature such as  [Chen et al.(2023)Chen, Han, and Zhang, Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen], we evaluated our method using the AUROC. Additionally, we report APs for anomaly detection and use AUPRO [Bergmann et al.(2020)Bergmann, Fauser, Sattlegger, and Steger] to assess pixel-level anomaly localization performance. Results are provided at both dataset-level and domain-level averages.
Table 1: Model performance on public benmarks. The top section reports image-level results (AUROC, AP), and the bottom section reports pixel-level results (AUROC, AUPRO). Best entries are marked with bold and second-best entries with underline.
Industrial Image-level Anomaly Detection (AUROC, AP) Dataset SAA CoCoOp WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP PILOT MVTec AD (64.2, 87.0) (88.2, 94.5) (91.9, 96.4) (86.5, 93.2) (91.6, 96.3) (91.3, 95.8) (92.1, 96.6) VisA (67.5, 76.2) (63.2, 68.5) (78.5, 81.4) (78.2, 81.3) (81.8, 85.1) (81.6, 84.1) (84.1, 85.7) BTAD (59.4, 89.2) (67.0, 77.6) (68.4, 71.1) (73.8, 68.9) (83.8, 88.4) (90.4, 90.7) (95.3, 96.8) SDD (68.9, 37.9) (75.1, 65.3) (84.5, 77.6) (80.0, 71.7) (85.5, 81.7) (81.4, 72.8) (87.9, 85.7) DAGM (87.3, 88.6) (87.7, 74.8) (91.9, 79.7) (94.6, 84.0) (96.1, 90.2) (97.3, 93.1) (98.5, 92.8) DTD-Synthetic (94.6, 93.7) (92.4, 94.8) (93.4, 92.8) (86.7, 95.2) (93.7, 97.3) (97.5, 98.2) (97.7, 98.4) Mean (73.7, 78.8) (78.9, 79.3) (84.8, 83.2) (83.3, 82.4) (88.8, 89.8) (89.9, 89.1) (92.6, 92.7) Medical Image-level Anomaly Detection (AUROC, AP) HeadCT (46.8, 68.0) (78.4, 78.8) (81.8, 80.2) (89.1, 89.4) (91.0, 92.3) (86.5, 81.3) (94.5, 96.0) BrainMRI (34.4, 76.7) (61.3, 44.9) (86.6, 91.5) (89.3, 90.9) (93.2, 94.7) (93.2, 89.7) (97.3, 97.9) Br35H (33.2, 67.3) (86.0, 87.5) (80.5, 82.2) (93.1, 92.9) (95.3, 96.0) (92.1, 90.9) (96.7, 97.7) Mean (38.1, 70.7) (75.2, 70.4) (83.0, 84.6) (90.5, 91.1) (93.2, 94.3) (90.6, 87.3) (96.2, 97.2) Industrial Pixel-level Anomaly Localization (AUROC, AUPRO) MVTec AD (75.5, 38.1) (44.4, 11.1) (85.1, 64.6) (87.6, 44.0) (90.7, 78.7) (89.4, 37.8) (90.3, 80.2) VisA (76.5, 31.6) (42.1, 12.2) (79.6, 56.8) (94.2, 86.8) (95.3, 85.1) (95.5, 77.8) (96.0, 87.2) BTAD (65.8, 14.8) (28.1, 6.5) (72.7, 27.3) (60.8, 25.0) (94.4, 73.6) (94.8, 32.5) (96.3, 74.9) SDD (78.8, 6.6) (24.4, 8.3) (68.8, 24.2) (79.8, 65.1) (90.7, 66.6) (71.7, 17.6) (92.4, 73.4) DAGM (62.7, 32.6) (17.5, 2.1) (87.6, 65.7) (82.4, 66.2) (95.6, 91.0) (91.1, 62.3) (96.0, 91.4) DTD-Synthetic (76.7, 60.6) (14.8, 3.0) (83.9, 57.8) (95.3, 86.9) (97.8, 91.1) (98.2, 86.2) (98.0, 90.2) Mean (72.7, 30.7) (28.6, 7.2) (79.6, 49.4) (83.4, 62.3) (94.1, 81.0) (90.1, 52.4) (94.8, 82.9)
Table 2: Comparison of PILOT vs. state-of-the-art on industrial benchmarks. $\mathbf{Img}$ and $\mathbf{Pix}$ denote image-level and pixel-level metrics, respectively. “w/o TTA” and “w/ TTA” indicate performance without and with TTA, respectively. Arrows indicate change in performance: $\uparrow$(increase >0.5>0.5), $\downarrow$(decrease <0.5<-0.5), and $\circ$(no significant change, $\pm 0.5$).
[…]
3.1 Comparison with State-of-the-Art
As reported in Table 1, PILOT establishes state-of-the-art performance across all benchmarks for both image-level anomaly detection and pixel-level localization in the zero-shot setting. This can be benefited from PILOT comprehensive approach in addressing critical challenges prevalent in ZSAD, that explicitly tackles domain shift in test phase. Prior approaches [Cao et al.(2023b)Cao, Xu, Sun, Cheng, Du, Gao, and Shen, Jeong et al.(2023)Jeong, Zou, Kim, Zhang, Ravichandran, and Dabeer] rely solely on predefined prompts, which may not generalize well to unseen domains, whereas methods such as [Chen et al.(2023)Chen, Han, and Zhang, Cao et al.(2024)Cao, Zhang, Frittoli, Cheng, Shen, and Boracchi, Zhou et al.(2022a)Zhou, Yang, Loy, and Liu, Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen] employ lightweight visual adapters or a single learnable prompt to capture anomaly cues, but often overfit to the auxiliary training set, limiting robustness under distribution shift. PILOT distinguishes itself by dynamically identifying the most relevant prompts for each input and refining only those learnable prompt parameters during inference using high-confidence pseudo-labels from the unlabeled target data. Note that while existing baselines can utilize the proposed image-level TTA strategies, these often degrade pixel-level localization performance and do not yield consistent improvements across both detection and localization tasks. On the contrary, PILOT uniquely achieves gains in both image-level and pixel-level metrics under TTA, a phenomenon analyzed in detail in Section 3.2. We also compare against state-of-the-art full-shot methods in Appendix F.1, which show that PILOT achieves image-level accuracy on par.
3.2 Effect of Test-Time Adaptation
Table 2 presents a comparison of all methods evaluated under a unified TTA protocol, where each method adapts learnable parameters across the test set using high-confidence pseudo-labels following PILOT’s label-free strategy. This fair setup reveals a key limitation of image-level TTA strategies in existing ZSAD models: while they can yield modest improvements in detection, they often destabilize pixel-level localization due to indiscriminate parameter updates guided by global objectives. This is particularly evident in methods like AnomalyCLIP, which use a single learnable prompt across all test images. In these cases, high-confidence pseudo-labels are often insufficient to represent the diversity of anomaly patterns in the target domain, resulting in imbalanced adaptation (see Appendix F.2) and degraded fine-grained localization. Thus, image-level TTA alone struggles to maintain the alignment between visual and semantic cues necessary for precise localization, especially in complex scenarios. AAnother significant limitation is conventional TTA methods’ susceptibility to noisy pseudo-labels, which propagate errors and degrade both detection and localization performance. PILOT’s adaptive prompt weighting addresses this by localizing the impact of label noise—updates concentrate on prompts closely aligned with each input, thus minimizing negative effects on the overall prompt pool.
3.3 Module Ablation
Table 3: Module ablation for PILOT on industrial datasets. SLP: Single Learnable Prompt; +Attr: SLP + Attribute Memory Bank; Full: SLP + Attr + Query & Prompt Pool (Full model). Metrics: Image = (AUROC, AP), Pixel = (AUROC, AUPRO).
Image-Level Pixel-Level
w/o TTA w/ TTA w/o TTA w/ TTA
[…]
Table 3 evaluates individual components of PILOT across industrial datasets, highlighting how our design stabilizes pixel-level predictions during TTA. Using a single prompt, similar to existing ZSAD methods, improves detection but severely degrades localization accuracy due to the lack of spatial supervision, causing visual-textual misalignment. Introducing the fixed attribute memory bank addresses this by providing stable semantic anchors, constraining model adaptation and preserving localization quality through parameter-free orthogonal projection. The complete PILOT framework integrates prompts from the learnable pool and attribute memory bank through a dual-branch fusion mechanism, dynamically weighted by adaptive query functions. During TTA, parameter updates are guided by alignment scores, ensuring prompts aligned with high-confidence samples undergo more substantial refinement. This adaptive update strategy mitigates negative transfer risks from unseen data, enhancing both image-level detection and pixel-level localization.
4 Conclusion
We have introduced PILOT, a ZSAD framework designed to address domain shifts by adaptively integrating anomaly cues through a dual-branch prompt aggregation mechanism, which combines a learnable prompt pool with semantic attribute embeddings. Our image-level TTA strategy further refines the prompt parameters using high-confidence pseudo-labels from unlabeled target data. Extensive experiments on 13 public benchmarks demonstrate that PILOT consistently achieves state-of-the-art performance in both anomaly detection and localization. Future work may explore advanced TTA strategies to jointly improve pixel-level localization and image-level detection, further enhancing ZSAD robustness.
Acknowledgements
The authors wish to acknowledge the financial support of McGill University, the Natural Sciences and Engineering Research Council of Canada (NSERC), and the Fonds de Recherche du Québec – Nature et technologies (FRQNT). We also acknowledge the computing resources provided by Calcul Québec and the Digital Research Alliance of Canada.
APPENDIX
This appendix provides supplementary details for the BMVC 2025 paper titled ”Zero-Shot Anomaly Detection with Dual-Branch Prompt Selection”.
  • Appendix A provides a comprehensive reviews of related works.
  • Appendix B reviews all datasets.
  • Appendix C provides information on the baseline methods.
  • Appendix D details implementation specifics.
  • Appendix E presents additional ablation studies for PILOT.
  • Appendix F presents a cross-dataset analysis of prompt diversity and semantics.
  • Appendix G reports medical localization results without TTA.
  • Appendix H provides a computational analysis of PILOT compared to existing baselines.
  • Appendix I compares AnovL with its TTA strategy.
  • Appendix J describes the construction of the attribute memory bank.
  • Appendix K provides the pseudo code for PILOT.
  • Appendix L shows qualitative visualizations of anomaly localization.
Appendix A Related Work
Zero-shot Anomaly Detection. ZSAD has gained increasing attention as a solution for identifying anomalies in unseen categories without requiring task-specific training data [Aota et al.(2023a)Aota, Tong, and Okatani, Jeong et al.(2023)Jeong, Zou, Kim, Zhang, Ravichandran, and Dabeer]. Leveraging the generalization capabilities of Vision–Language Models (VLMs) such as CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.], many approaches extend these models by linking image features to textual descriptions of normal and anomalous states. For instance, CLIP-AD [Chen et al.(2024)Chen, Zhang, Tian, He, Zhang, Wang, Wang, and Liu] and ZOC [Esmaeilpour et al.(2022)Esmaeilpour, Liu, Robertson, and Shu] are early studies that leverage CLIP for ZSAD, but primarily focus on anomaly classification rather than localization. WinCLIP [Jeong et al.(2023)Jeong, Zou, Kim, Zhang, Ravichandran, and Dabeer] uses multiple handcrafted prompts with iterative patch-based forward passes, and April-GAN [Chen et al.(2023)Chen, Han, and Zhang] refines local semantics via fine-tuning on an auxiliary dataset; both methods address image-level anomaly detection and pixel-level localization. More recent prompt learning methods, such as AnomalyCLIP [Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen], AnomalyGPT [Gu et al.(2024)Gu, Zhu, Zhu, Chen, Tang, and Wang], and AdaCLIP [Cao et al.(2024)Cao, Zhang, Frittoli, Cheng, Shen, and Boracchi], optimize object-agnostic text prompts in a single pass using auxiliary dataset. However, reliance on auxiliary datasets may introduce bias and compromise the intended zero-shot setting. This trade-off has been examined by Marzullo et al. [Marzullo and Ranzini(2024)], who observed that while avoiding auxiliary data reduces fine-tuning overhead, static prompts often fail to adapt to domain shifts; conversely, fine-tuning on auxiliary sets can improve accuracy but risks dataset bias and reduced generalization. Our work aims at addressing these challenges by introducing a dual-branch prompt learning mechanism that combines learnable and fixed prompts, enabling more robust and adaptive anomaly detection under domain shift.
CLIP Adaptation. Recent advances in CLIP fine-tuning primarily rely on prompt learning, in which learnable tokens are introduced into the text encoder to align visual features with related semantics [Zhou et al.(2022b)Zhou, Yang, Loy, and Liu]. Approaches such as CoCoOp [Zhou et al.(2022a)Zhou, Yang, Loy, and Liu] and visual prompt tuning [Jia et al.(2022)Jia, Tang, Chen, Cardie, Belongie, Hariharan, and Lim] condition the prompts on image content to improve robustness under domain shift. MaPLe [Khattak et al.(2023)Khattak, Rasheed, Maaz, Khan, and Khan] extends prompt learning by inserting learnable tokens into both the visual and text encoders of CLIP, jointly optimizing them to enhance cross-modal alignment. These techniques achieve parameter-efficient adaptation by updating only the prompt representations and, in some cases, only CLIP’s final visual projection layer [Fahes et al.(2024)Fahes, Vu, Bursuc, Pérez, and de Charette]. Following this rationale, our method fine-tunes the visual projection layer to minimize additional parameters.
Test Time Adaptation. A related direction for adapting CLIP-based models, particularly under domain shift, is TTA. In TTA, the model updates its parameters using unlabeled test data, often guided by pseudo-labels or self-supervised objectives. For instance, self-supervised losses based on the model’s own outputs can fine-tune representations without ground-truth labels [Sun et al.(2020)Sun, Wang, Liu, Miller, Efros, and Hardt], and momentum-based pseudo-labeling helps enforce consistency across different augmented views [Goyal et al.(2022)Goyal, Sun, Raghunathan, and Kolter]. Other TTA strategies include entropy minimization, which guides the model to assign higher confidence to its predictions on unlabeled data, and the online updating of batch normalization statistics, which helps the model better align with the target distribution during test time. These approaches have been shown to improve robustness and generalization in image classification [Nado et al.(2020)Nado, Padhy, Sculley, D’Amour, Lakshminarayanan, and Snoek] and segmentation [Colomer et al.(2023)Colomer, Dovesi, Panagiotakopoulos, Carvalho, Härenstam-Nielsen, Azizpour, Kjellström, Cremers, and Poggi]. However, recent studies have found that these techniques are often less effective when the domain shift is large or when the model relies on strong vision-language priors, as in the case of CLIP. To our knowledge, AnovL [Deng et al.(2023)Deng, Zhang, Bao, and Li] is the only TTA method for zero-shot anomaly localization, which employs an adapter module for pixel-level refinement and thus introduces additional parameters while focusing solely on localization. In contrast, our method adaptively updates the relevant learnable prompts, incurring no additional parameters and delivering consistent improvements in both image-level anomaly detection and pixel-level anomaly localization.
Appendix B Dataset
B.1 Benchmark Datasets
Appendix B provides a comprehensive review of the thirteen benchmark datasets utilized in our study, as summarized in Table 4. To rigorously evaluate both image-level detection and pixel-level localization capabilities of PILOT, we select six industrial inspection datasets. DAGM [Wieler and Hahn(2007)] and DTD-Synthetic [Aota et al.(2023b)Aota, Tong, and Okatani] comprise textured surfaces exhibiting subtle variations in material appearance, presenting unique challenges for texture-based anomaly detection due to the fine-grained and ambiguous nature of defects. In contrast, MVTec AD [Bergmann et al.(2019)Bergmann, Fauser, Sattlegger, and Steger], VisA [Zou et al.(2022)Zou, Jeong, Pemula, Zhang, and Dabeer], BTAD [Mishra et al.(2021)Mishra, Verk, Fornasier, Piciarelli, and Foresti], and SDD [Tabernik et al.(2020)Tabernik, Sela, Skvarc, and Skocaj] focus on object-centric industrial scenarios, with defects such as cracks, scratches, dents, and missing components distributed across everyday manufactured and synthetic objects. These datasets are characterized by diverse object categories and high variability in anomaly manifestation, supporting comprehensive benchmarking of localization and generalization performance.
Table 4: Summary of the 13 public benchmark datasets. “Ind” = Industrial, “Med” = Medical, “Def” = Defect detection, “AD” = Anomaly detection, “AL” = Anomaly localization.
Dataset Cls. Norm Anom Total Dom. Task
MVTec AD 15 467 1258 1725 Ind Def
VisA 12 962 1200 2162 Ind Def
BTAD 3 451 290 741 Ind Def
SDD 1 181 74 255 Ind Def
DAGM 10 6996 1054 8050 Ind Def
DTD-Synthetic 12 357 947 1304 Ind Def
HeadCT 1 100 100 200 Med AD
BrainMRI 1 98 155 253 Med AD
Br35H 1 1500 1500 3000 Med AD
ISIC 1 0 379 379 Med AL
CVC-ClinicDB 1 0 612 612 Med AL
CVC-ColonDB 1 0 380 380 Med AL
Kvasir 1 0 1000 1000 Med AL
To assess clinical applicability, we include seven medical imaging benchmarks. Three brain scan datasets, HeadCT [Kitamura(2018)], BrainMRI [Kanade and Gumaste(2015)], and Br35H [Hamada(2020)], comprise both healthy and pathological subjects, facilitating binary anomaly classification tasks on volumetric neuroimaging data. These datasets encompass a spectrum of anatomical and pathological variations, thereby enabling evaluation of robustness to inter-patient heterogeneity and rare conditions. The remaining four datasets provide pixel-wise lesion annotations and consist exclusively of abnormal samples. ISIC [Codella et al.(2018)Codella, Gutman, Celebi, Helba, Marchetti, Dusza, Kalloo, Liopyris, Mishra, Kittler, et al.] features dermoscopic images annotated for skin lesions, while CVC-ClinicDB [Bernal et al.(2015)Bernal, Sánchez, Fernández-Esparrach, Gil, Rodríguez, and Vilariño], CVC-ColonDB [Tajbakhsh et al.(2015)Tajbakhsh, Gurudu, and Liang], and Kvasir [Jha et al.(2020)Jha, Smedsrud, Riegler, Halvorsen, De Lange, Johansen, and Johansen] contain endoscopic images of the gastrointestinal tract, highlighting a range of polyp and ulcer appearances. Since these medical localization datasets lack normal examples, test-time adaptation is not performed for them; their localization results without TTA are reported in Appendix G.
B.2 Domain Discrepancy Analysis
To validate that MVTec AD and VisA indeed represent statistically distinct domains suitable for strict zero-shot evaluation, we quantitatively assessed the distributional discrepancy between their CLIP global feature embeddings. Specifically, we employed Maximum Mean Discrepancy (MMD2[Gretton et al.(2012)Gretton, Borgwardt, Rasch, Schölkopf, and Smola], Fréchet Inception Distance (FID) [Heusel et al.(2017)Heusel, Ramsauer, Unterthiner, Nessler, and Hochreiter], and the Kolmogorov–Smirnov (KS) [Massey Jr(1951)] test. Our analysis yielded $\mathrm{MMD}^{2}=0.0328$ ($p$-value =0.0099=0.0099), indicating a significant difference in embedding distributions. The FID score between the two datasets is $51.9$, further corroborating a substantial distributional gap. Additionally, the KS test computed across all feature dimensions revealed that the minimum, mean, and maximum $p$-values are 0, $5.16\times 10^{-5}$, and $0.0329$, respectively, suggesting statistically significant differences in nearly every dimension of the embedding space. These results collectively confirm that MVTec AD and VisA are non-overlapping in the learned representation space, thus justifying their use as auxiliary and target datasets in our zero-shot anomaly detection experiments.
Appendix C Baseline
In Appendix C, we compare PILOT against six representative zero-shot anomaly detection methods, summarized as follows:
  • SAA [Cao et al.(2023b)Cao, Xu, Sun, Cheng, Du, Gao, and Shen]: Relies on a pretrained Grounding DINO to propose candidate anomaly regions, which are then segmented by the Segment Anything Model (SAM) without any additional training.
  • CoCoOp [Zhou et al.(2022a)Zhou, Yang, Loy, and Liu]: Replaces fixed CLIP prompt tokens with learnable context vectors while keeping the backbone frozen. Offers both a shared “unified” prompt and class-specific prompts to adapt flexibly to unseen categories.
  • WinCLIP [Jeong et al.(2023)Jeong, Zou, Kim, Zhang, Ravichandran, and Dabeer]: Splits the image into overlapping windows, encodes each via CLIP to produce localized anomaly scores, and aggregates window predictions through an ensemble of token embeddings—improving spatial localization without fine-tuning.
  • APRIL-GAN [Chen et al.(2023)Chen, Han, and Zhang]: Builds on WinCLIP by refining text prompts for tighter vision–language alignment and injecting lightweight learnable projections into the visual encoder, thereby sharpening anomaly masks while retaining zero-shot generality.
  • AnomalyCLIP [Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen]: Learns object-agnostic “normal” and “anomalous” text prompts, then employs a dense patch-to-prompt attention module (DPAM) to align each visual patch embedding with the appropriate prompt for pixel-level localization.
  • AdaCLIP [Cao et al.(2024)Cao, Zhang, Frittoli, Cheng, Shen, and Boracchi]: Combines static prompts shared across all inputs for coarse adaptation with dynamic, image-conditioned prompts for fine-grained adjustment. Both prompt sets are jointly learned on auxiliary anomaly data to balance detection accuracy and localization precision.
Appendix D Implementation Details
D.1 PILOT Details
Appendix D provides full implementation details of PILOT. We implement PILOT atop the open-source CLIP ViT-L/14@336px backbone, freezing all transformer weights except for the final visual-to-joint projection layer, which is jointly optimized alongside our prompt modules. Our learnable prompt pool $\mathcal{P}$ contains $K=5$ paired prompts that each comprise $L=12$ learnable vectors for “normal” and “anomalous” semantics, while a fixed attribute memory bank (see Appendix J) provides semantic tokens extracted from structured templates. All images are uniformly resized to $518\times 518$ pixels, and we extract both the [CLS] token $x_{\mathrm{cls}}$ and patch embeddings $\{X^{(M)}\}$ at layers 6, 12, 18, and 24 of the CLIP encoder. All prompt embeddings are initialized by sampling from a Gaussian distribution $\mathcal{N}(0,\,0.02)$. PILOT is trained for five epochs using AdamW with a base learning rate of $10^{-3}$, a batch size of 32. During test-time adaptation, we fix the pseudo-label fraction at $\rho=0.25$. All experiments are conducted in PyTorch 2.1.0 on a single NVIDIA Ada A5000 GPU with 32 GB of memory.
D.2 Anomaly State Descriptors
The following anomaly state descriptors are used in our nearest-neighbor analysis, as described in Appendix F of the main text. These descriptors were selected to cover a wide range of mechanical, surface, electrical, pressure-related, assembly, and contamination defects:
  • Mechanical damage: damaged, scratched, cracked, chipped, broken, bent, punctured, fractured, delaminated, weld defect, loose component, misaligned, deformed, twisted, buckled
  • Surface & corrosion: rusty, corroded, oxidized, pitted, eroded, paint peeling, stained, contaminated, oil leak, grease smear
  • Thermal & electrical: overheated, burnt, scorched, melted, short-circuited, electrical arcing, voltage spike damage
  • Sealing & pressure: leaking, sealing failure, gasket blown, under-pressure, over-pressure, gas leak
  • Wear & fatigue: worn bearing, surface fatigue crack, abrasion mark, material fatigue, delamination
  • Assembly & alignment: loose screw, missing bolt, stripped thread, cross-threaded, flange misfit
  • Contamination & blockages: clogged, blocked channel, foreign particle inclusion, scale buildup, slag inclusion
  • Other: vibration damage, impact dent, surface blistering, vacuum leak, structural crack
D.3 Entmax1.5 Details
Let $\boldsymbol{\kappa}=(\kappa_{1},\dots,\kappa_{D})\in\mathbb{R}^{D}$ be a vector of input scores, where each $\kappa_{i}\in\mathbb{R}$ denotes the unnormalized score for the $i$-th dimension of interest, and write the probability simplex as
$\Delta^{D}=\bigl{\{}\boldsymbol{\chi}\in\mathbb{R}^{D}\mid\chi_{i}\geq 0,\;\sum_{i=1}^{D}\chi_{i}=1\bigr{\}},$ (9)
where $\Delta^{D}$ is the set of all $D$-dimensional probability distributions.
The $\mathrm{entmax}_{1.5}$ operator produces a sparse probability vector $\boldsymbol{\chi}\in\Delta^{D}$ via
$\chi_{i}=\bigl{[}\max\bigl{(}0,\;\kappa_{i}-\tau(\boldsymbol{\kappa})\bigr{)}\bigr{]}^{2},\quad i=1,\dots,D,$ (10)
where the scalar $\tau(\boldsymbol{\kappa})\in\mathbb{R}$ is chosen so that $\sum_{i=1}^{D}\chi_{i}=1$. Intuitively, any entry with $\kappa_{i}\leq\tau(\boldsymbol{\kappa})$ is set to zero, and the remaining values are shifted, squared, and renormalized. This yields a differentiable, piecewise-quadratic mapping that retains only the most salient dimensions in its support. In PILOT, injecting Gumbel noise before applying entmax1.5 ensures that only the highest-scoring attribute cues survive to influence the final text embedding, providing robust, sparse weighting during training.
Appendix E More Ablation Study
E.1 Comparison with SOTA full-shot method
Table 5: Performance comparison on industrial anomaly detection benchmarks. PILOT is evaluated in the zero-shot setting; all other methods are full-shot. Top section: image-level (AUROC, AP); bottom: pixel-level (AUROC, AUPRO). Best results in bold, second-best underlined.
Industrial Image-level Anomaly Detection (AUROC, AP)
Dataset PILOT PatchCore CDO RD4AD
MVTec AD (92.1, 96.6) (99.0, 99.7) (97.1, 97.2) (98.7, 99.4)
VisA (84.1, 85.7) (94.6, 95.9) (95.0, 95.1) (95.3, 95.7)
BTAD (95.3, 96.8) (93.2, 98.6) (97.6, 97.7) (93.8, 96.8)
SDD (87.9, 85.7) (91.8, 91.6) (96.0, 96.2) (86.8, 81.3)
DAGM (98.5, 92.8) (95.4, 90.3) (95.1, 95.0) (92.9, 79.1)
DTD-Synthetic (97.7, 98.4) (97.5, 97.3) (96.8, 96.9) (96.2, 96.3)
Industrial Pixel-level Anomaly Localization (AUROC, AUPRO)
Dataset PILOT PatchCore CDO RD4AD
MVTec AD (90.3, 80.2) (98.1, 92.8) (98.2, 98.1) (97.8, 97.6)
VisA (96.0, 87.2) (98.5, 92.2) (99.0, 99.2) (98.4, 98.9)
BTAD (96.3, 74.9) (97.4, 74.4) (98.1, 98.2) (97.5, 97.8)
SDD (92.4, 73.4) (87.9, 46.3) (97.9, 97.7) (92.2, 92.0)
DAGM (96.0, 91.4) (95.9, 87.9) (97.3, 97.2) (96.8, 96.9)
DTD-Synthetic (98.0, 90.2) (98.2, 98.1) (98.3, 98.5) (97.6, 97.7)
In this comparison, we evaluate PILOT alongside three state-of-the-art fully supervised methods—PatchCore [Roth et al.(2022)Roth, Pemula, Zepeda, Schölkopf, Brox, and Gehler], CDO [Cao et al.(2023a)Cao, Xu, Liu, and Shen], and RD4AD [Deng and Li(2022)]—across six widely used industrial benchmarks. Here, a full-shot method refers to a model trained with full access to samples from the target domain. While these baselines leverage direct access to target domain knowledge, PILOT operates solely with pretrained vision–language representations and requires no information from the target domain. Remarkably, PILOT achieves detection accuracy on par with these full-shot models, even when faced with unseen textures or novel defect types. It matches PatchCore’s memory-based robustness, rivals the collaborative feature learning of CDO, and maintains high sensitivity to both global and local irregularities, comparable to RD4AD’s reconstruction-driven approach. However, there remains room for improvement in pixel-level localization accuracy, which we identify as a promising direction for future research. These results underscore the significant potential of zero-shot anomaly detection frameworks in real-world industrial scenarios.
E.2 Pseudo label Statistic
Table 6: TTA Dataset selection distribution statistics across datasets under different pseudo-labeling ratios $\rho$. M1: Shannon Evenness, M2: Coefficient of Variation, M3: Percentage of Noisy Label.
Dataset Class $\rho=$5% $\rho=$15% $\rho=$25% $\rho=$35% $\rho=$50% M1 M2 M3 M1 M2 M3 M1 M2 M3 M1 M2 M3 M1 M2 M3 MVTec AD 15 0.76 1.21 10.3% 0.97 0.45 8.1% 0.93 0.71 7.8% 0.95 0.59 11.2% 1.00 0.00 27.3% VisA 12 0.79 0.97 12.0% 0.98 0.27 13.2% 0.98 0.35 14.2% 0.99 0.31 16.1% 1.00 0.00 31.4% BTAD 3 0.79 0.76 21.6% 0.86 0.70 12.6% 0.93 0.50 17.3% 0.95 0.39 20.4% 1.00 0.00 36.5% SDD 1 0.72 0.59 0.0% 0.61 0.69 7.6% 0.63 0.68 6.3% 0.74 0.58 10.1% 1.00 0.00 39.6% DAGM 10 0.73 1.35 3.9% 0.88 0.76 1.3% 0.95 0.56 0.7% 0.97 0.43 0.9% 1.00 0.00 6.9% DTD-Synthetic 12 0.72 1.34 3.0% 0.96 0.47 6.1% 0.95 0.53 4.6% 0.96 0.48 5.2% 1.00 0.00 20.7%
In Table 6 we report three statistics that describe how our TTA selection (See Section 2.5 of the main paper) distributes pseudo-labels over the target dataset $D_{\mathrm{tar}}$. Let $B=\lvert D_{\mathrm{tar}}\rvert$ and recall that we form the top-$\rho$ set $A$ and bottom-$\rho$ set $N$, with $S=A\cup N$. For each class $i$ in the set of classes $\mathcal{I}$, we write $g_{i}$ for the fraction of samples in $S$ belonging to class $i$.
  • Shannon evenness $E$ is defined as
    $E=-\frac{\sum_{i\in\mathcal{I}}g_{i}\ln g_{i}}{\ln\lvert\mathcal{I}\rvert}$.
    Values near 1 indicate that the selected samples $S$ are spread uniformly across all classes. This is referred as M1 in Table 6.
  • Coefficient of variation $\mathrm{CV}$ is
    $\mathrm{CV}=\frac{\sigma(g)}{\mu(g)},$
    where $\mu(g)$ and $\sigma(g)$ are the mean and standard deviation of $\{g_{i}\}$. A low $\mathrm{CV}$ signals a balanced class-wise count. This is referred as M2 in Table 6.
  • Noisy label rate is the fraction of indices in $S$ whose pseudo-label $\tilde{y}_{j}$ disagrees with the ground-truth label $y_{j}$. This is referred as M3 in Table 6.
As $\rho$ grows, we observe that $E$ increases and $\mathrm{CV}$ decreases, confirming that selecting a larger fraction of samples both diversifies class coverage and balances the counts. At the same time, the noisy label rate rises, reflecting the classic trade-off: higher coverage under TTA comes at the expense of pseudo-label accuracy.
E.3 Other TTA Method
To better understand the impact of different TTA strategies on anomaly detection performance, we conduct an ablation study comparing several representative TTA methods. As shown in Table 7, we evaluate TENT [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell], BN Adaptation [Lim et al.(2023)Lim, Kim, Choo, and Choi], and soft pseudo labeling alongside our proposed PILOT framework.
Table 7: Ablation of different TTA methods on industrial anomaly detection datasets. Metrics reported: Image-level (AUROC, AP) and Pixel-level (AUROC, AUPRO).
TTA Method Image-level Pixel-level
AUROC AP AUROC AUPRO
TENT 74.3 73.5 61.6 40.9
BN Adaptation 70.5 69.8 55.2 34.5
Soft Pseudo Label 91.0 91.2 92.0 79.8
PILOT 92.6 92.7 94.8 82.9
Table 7 presents the results of different TTA strategies on industrial anomaly detection benchmarks. Both TENT and BN Adaptation exhibit weak performance in this setting, particularly for pixel-level localization, suggesting that these conventional TTA methods are not well suited for adapting CLIP-based models. Recent literature has also observed that methods such as TENT, which adapt only batch normalization parameters, often struggle when applied to models with large distribution shifts or strong vision-language priors, as is the case with CLIP [Liang et al.(2025)Liang, He, and Tan]. In contrast, the soft pseudo labeling strategy leads to more stable improvements, although its cautious approach may be less aggressive in leveraging adaptation signals. Our findings suggest that confidence-based pseudo-labeling, as used in the PILOT TTA framework, provides the most effective adaptation and yields the best overall performance in this scenario.
E.4 Effect of Learnable Prompt Pool Size
Table 8: Effect of prompt pool size $K$ on the industrial benchmark. Best in each row is bold.
Level Metric $K$
2 3 4 5 7 10 16 32 64
Image-level AUROC 85.9 88.6 91.3 92.6 92.4 92.4 92.9 93.2 91.7
AP 87.2 90.8 91.9 92.7 92.0 92.0 93.5 93.8 92.2
Pixel-level AUROC 94.1 93.7 94.6 94.8 91.7 88.0 89.4 87.1 83.7
AUPRO 82.1 81.0 82.4 82.9 80.9 77.2 78.3 76.3 73.8
Table 8 examines how changing the number of learnable prompts influences performance. As the pool size increases, image-level metrics rise steadily and then level off once a moderate number of prompts is reached, indicating improved anomaly discrimination with greater prompt diversity. However, pixel-level metrics follow a different trend: they increase up to a moderate pool size but then decline with further increases, reflecting reduced generalization for localization. This divergence shows that while a larger pool improves global detection, overly large pools lead to redundancy and overfitting at the pixel level, making a moderate pool size optimal for balancing both objectives.
E.5 Effect of Attribute Memory Bank Size.
Table 9: Effect of attribute-memory bank size $C$ on the industrial benchmark. Best in each row is bold.
Level Metric $C$
5 25 50 75 100 150
Image-level AUROC 88.1 92.1 92.5 92.6 92.7 92.2
AP 88.4 92.0 92.4 92.7 92.6 92.3
Pixel-level AUROC 87.7 93.9 94.5 94.8 94.6 94.4
AUPRO 76.1 82.2 82.7 82.9 82.8 82.5
Table 9 shows the impact of varying the number of attribute templates. Expanding the memory bank from a small to a moderate size leads to clear improvements in both image-level and pixel-level performance, as the model benefits from richer semantic information. Beyond this range, image-level metrics stabilize, while pixel-level gains plateau and may slightly decrease with the largest banks. This suggests that adding more attributes beyond a certain point does not further improve generalization, and that a compact but diverse attribute bank is sufficient to provide strong semantic guidance.
E.6 Effect of $\rho$ fraction
Table 10: Effect of pseudo-label fraction $\rho$ (%) during TTA (industrial datasets). Best value in each column is bold.
Level Metric $\rho$ (%)
5 15 25 35 50
Image-level AUROC 92.4 93.0 92.6 90.1 88.5
AP 92.5 93.1 92.7 90.2 88.7
Pixel-level AUROC 92.4 93.6 94.8 92.7 91.5
AUPRO 81.0 81.9 82.9 79.9 78.0
We vary the pseudo-label fraction $\rho$ during TTA, as shown in Table 10. Using a small set of high-confidence pseudo labels leads to improvements in image-level performance compared to no adaptation, as the adaptation process focuses on the most reliable samples. Notably, pixel-level performance remains largely unchanged when only a small subset is used, but begins to improve as $\rho$ increases to an intermediate fraction. At this optimal point, the model is able to leverage a greater diversity of anomaly patterns without introducing substantial label noise. Beyond this, further increases in $\rho$ admit less reliable pseudo labels, resulting in a sharp decline in both detection and localization metrics.
E.7 Prompt Aggregation Ablation
Table 11: Ablation for prompt aggregation. “w/o TTA” and “w/ TTA” indicate performance without and with TTA, respectively. “Extra” indicates whether additional trainable parameters are used. OP: orthogonal projector as used in PILOT; GF: Gate Fusion [Arevalo et al.(2017)Arevalo, Solorio, Montes-y Gómez, and González]; CA: Cross Attention [Chen et al.(2021)Chen, Fan, and Panda].
[…]
Table 11 compares three prompt aggregation strategies and reveals that methods like Gate Fusion [Arevalo et al.(2017)Arevalo, Solorio, Montes-y Gómez, and González] and Cross Attention [Chen et al.(2021)Chen, Fan, and Panda], while capable of improving image-level anomaly detection, tend to destabilize pixel-level localization during TTA. While these approaches may leverage the attribute memory bank as a semantic anchor, their extra parameters are directly updated using noisy pseudo-labels and only image-level supervision. This makes them prone to parameter drift, which can erode the stability needed for precise localization under domain shift. In contrast, the orthogonal projector used in PILOT achieves reliable gains across both detection and localization tasks without introducing extra parameters, as it keeps the attribute embedding fixed as a semantic reference and allows only the prompt representation to adapt. This encourages adaptation in directions that are truly complementary, while avoiding redundancy and overfitting.
E.8 Effect of Orthogonal Projection
Figure 2: Histogram of normalized projection residuals for normal and anomaly prompts (see Eq. 5). The y-axis shows the fraction of samples per bin. Higher values indicate greater orthogonality to the anchor, while lower values reflect stronger alignment.
To evaluate the effectiveness of our orthogonal projection strategy for prompt aggregation (Equation 5), we analyze the distribution of normalized projection residuals defined as:
$\frac{\|T_{(\cdot)}-\mathrm{Proj}_{T_{\mathrm{att}}}(T_{(\cdot)})\|_{2}}{\|T_{(\cdot)}\|_{2}}$ (11)
for both learnable prompts $T_{n}$ (normal) and $T_{a}$ (anomaly), as shown in Figure 2. The resulting histograms exhibit consistently nonzero residuals, which are clearly separated between normal and anomalous samples. This indicates that the learnable prompts encode substantial information complementary to the attribute embedding $T_{\mathrm{att}}$, rather than merely duplicating its semantics. Notably, anomalous prompts display larger orthogonal components, highlighting the model’s ability to leverage diverse information captured by the learnable prompts. Empirically, these findings confirm that the orthogonal projection mechanism extracts and integrates useful, non-redundant cues into the fused embedding, thereby validating our parameter-free aggregation approach.
Appendix F Prompt Diversity
To examine the diversity within our learnable prompt pool $\mathcal{P}$, we analyzed pairwise relationships among the learned prompt embeddings across multiple industrial datasets. Specifically, we measured the pairwise cosine similarity between prompt embeddings within each dataset. Figure 3(a) summarizes these findings using boxplots. The results demonstrate consistent diversity across all evaluated datasets, with median similarity values generally below 0.5, indicating that the learned prompts capture distinct representations.
(a) Prompt-pair similarity
(b) Prompt contribution
Figure 3: (a) Prompt-pair similarity distributions: each horizontal boxplot shows absolute cosine similarity between learned prompt embeddings. Values close to 0 indicate high diversity. (b) Prompt contribution: normalized contribution of a learnable prompt from the pool $\mathcal{P}$, aggregated over test images. Higher values indicate greater involvement during TTA.
To further understand how this diversity is practically leveraged, Figure 3(b) illustrates the relative contributions of each prompt in the pool across multiple datasets. The observed variations suggest that rather than relying on a single dominant prompt, the model adaptively emphasizes different prompts based on specific dataset characteristics. This adaptive utilization indicates that the prompt pool effectively leverages its inherent diversity to accommodate the varying requirements of different anomaly detection scenarios.
Table 12: Top-3 anomaly state matches for each learned prompt on the VisA dataset, retrieved via FAISS nearest neighbor search between the prompt embedding and the CLIP text embedding of “a photo of a [anomaly state] object”. For each prompt, $\circ$ indicates matches after finetuning on the auxiliary set ($\mathcal{D}_{\mathrm{aux}}$), and $\bullet$ indicates matches after TTA on the target set ($\mathcal{D}_{\mathrm{tar}}$). The top-3 matches are listed in descending order of cosine similarity.
Prompt Top-3 Matches
1 $\circ$ misaligned, bent, over-pressure
$\bullet$ bent, misaligned, flange misfit
2 $\circ$ abrasion mark, delamination, oxidized
$\bullet$ bent, flange misfit, clogged
3 $\circ$ damaged, corroded, weld defect
$\bullet$ damaged, impact dent, eroded
4 $\circ$ damaged, cracked, weld defect
$\bullet$ damaged, voltage spike damage, weld defect
5 $\circ$ damaged, paint peeling, weld defect
$\bullet$ damaged, voltage spike damage, contaminated
Having established the inherent diversity within our learnable prompt pool $\mathcal{P}$, we further investigated how this diversity corresponds to meaningful anomaly characteristics expressed in natural language. To accomplish this, we conducted a nearest-neighbor analysis using FAISS [Douze et al.(2024)Douze, Guzhva, Deng, Johnson, Szilvasy, Mazaré, Lomeli, Hosseini, and Jégou], comparing anomaly prompt embeddings ($T_{a}$) to embeddings generated from textual templates of the form “a photo of a [anomaly state] object.” Various anomaly state descriptors were used to represent potential real-world anomalies.
Table 12 lists the top-3 nearest textual descriptors for each prompt embedding, both before and after TTA. Additional results are provided in Appendix G. Prompts were initially trained on the MVTec auxiliary set and subsequently adapted to the VisA target set. The nearest-neighbor analysis revealed clear and meaningful alignments between the learned prompts and natural language anomaly descriptors. Prompts originally associated with anomaly states such as “misaligned,” “bent,” or “eroded” consistently maintained these semantic associations after adaptation, showing subtle shifts that reflect nuances in the target dataset. These observations underscore that the diversity within our learned prompt pool corresponds meaningfully to recognizable natural-language concepts.
Table 13: Top-3 anomaly state matches for each learned prompt (1155) on six datasets. Ranks are indicated by ①, ②, and ③. Each match is the closest to the template “a photo of a [anomaly state] object”.
Prompt MVTec AD VisA BTAD SDD DAGM DTD-Synthetic
1 ① rusty
② abrasion mark
③ flange misfit
① bent
② misaligned
③ flange misfit
① overheated
② short-circuited
③ misaligned
① bent
② misaligned
③ electrical arcing
① bent
② eroded
③ electrical arcing
① delamination
② under-pressure
③ over-pressure
2 ① punctured
② electrical arcing
③ short-circuited
① bent
② flange misfit
③ clogged
① abrasion mark
② weld defect
③ surface blistering
① abrasion mark
② chipped
③ weld defect
① abrasion mark
② grease smear
③ chipped
① electrical arcing
② delamination
③ under-pressure
3 ① damaged
② impact dent
③ cracked
① damaged
② impact dent
③ eroded
① damaged
② weld defect
③ abrasion mark
① damaged
② abrasion mark
③ scratched
① surface fatigue crack
② grease smear
③ scratched
① damaged
② short-circuited
③ eroded
4 ① bent
② gasket blown
③ cross-threaded
① damaged
② voltage spike damage
③ weld defect
① damaged
② cracked
③ broken
① weld defect
② surface fatigue crack
③ damaged
① surface fatigue crack
② cracked
③ damaged
① cracked
② damaged
③ broken
5 ① scratched
② grease smear
③ overheated
① damaged
② voltage spike damage
③ contaminated
① damaged
② weld defect
③ abrasion mark
① weld defect
② damaged
③ surface fatigue crack
① surface fatigue crack
② structural crack
③ weld defect
① surface fatigue crack
② grease smear
③ structural crack
Table 13 further extends this semantic exploration, reporting the top-3 nearest textual matches for each learned prompt (1155) across all six industrial anomaly detection datasets. Each match is derived from the template "a photo of a [anomaly state] object."
This broader analysis underscores that individual prompts exhibit varying associations with different anomaly states depending on the dataset. Such variability demonstrates the model’s capacity to adapt prompt representations across diverse domains. While some semantic associations remain stable, others shift considerably, reflecting both shared and domain-specific semantic cues. Collectively, these findings reinforce the effectiveness of our prompt pool in capturing diverse anomaly characteristics and adapting its semantic emphasis in response to domain-specific features.
Appendix G Medical anomaly localization results
Table 14: Comparison of model performance on medical anomaly localization datasets. Pixel-level results (AUROC, AUPRO) in parentheses. Best entries are marked with bold and second-best entries with underline.
Medical Pixel-level Anomaly Localization (AUROC, AUPRO) Dataset SAA CoCoOp WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP PILOT w/o TTA ISIC (83.8, 74.2) (51.7, 15.9) (83.3, 55.1) (89.4, 77.2) (87.4, 74.5) (84.4, 54.5) (87.8, 74.2) CVC-ColonDB (71.8, 31.5) (40.5, 2.6) (70.3, 32.5) (78.4, 64.6) (88.5, 79.3) (88.0, 63.9) (88.6, 78.2) CVC-ClinicDB (66.2, 29.1) (34.8, 2.4) (51.2, 13.8) (80.5, 60.7) (93.0, 81.6) (94.4, 73.5) (92.1, 78.5) Kvasir (45.9, 13.3) (44.1, 3.5) (69.7, 24.5) (75.0, 36.2) (93.2, 59.9) (94.6, 26.2) (94.0, 58.8) Mean (66.9, 37.0) (42.8, 6.1) (68.6, 31.5) (80.8, 59.7) (90.5, 73.8) (90.4, 54.5) (90.6, 72.4)
Medical anomaly localization results are reported in Table 14, with all evaluations conducted without TTA. Since medical datasets lack explicit “normal” examples, our TTA strategy cannot be applied; nevertheless, PILOT still produces competitive localization maps, achieving the highest mean AUROC and the second-best AUPRO across benchmarks. These results underscore the generality of our dual-branch prompt learning framework: even in the absence of adaptation, dynamically weighting prompts with the most relevant anomaly cues deliver state-of-the-art performance in medical domains.
Appendix H Computational Analysis
Table 15: Comparison of computational resource requirements for PILOT and main competitors. GFLOPs, latency, and GPU memory are reported for the test phase. Trainable parameters are reported for the training phase in all methods; for PILOT, test-time adaptation parameters are also shown (train / test).
Method Params (M) GFLOPs Latency (ms) GPU Mem (MB) AnomalyCLIP 5.56 482.06 107.4 2051.7 AdaCLIP 10.67 2872.37 219.0 8752.5 PILOT 6.40 (Train) / 5.61 (Test) 491.39 109.9 2398.4
Table 15 compares PILOT against our main competitors, AnomalyCLIP and AdaCLIP, in terms of model size, computational cost, inference speed, and GPU memory footprint. All measurements were obtained on an NVIDIA Ada A5000 GPU, as detailed in Appendix D.1. PILOT exhibits a resource profile closely aligned with AnomalyCLIP, and significantly outperforms AdaCLIP across all measured metrics. Importantly, PILOT employs a two-phase process: during the initial training phase, it utilizes a moderately sized model, while in the TTA phase, only the learnable prompt pool ($\mathcal{P}$) is updated. By limiting parameter updates to $\mathcal{P}$ during deployment, PILOT further reduces computational requirements, enhancing its practical suitability for near real-time anomaly detection in resource-constrained environments.
Appendix I Compared with AnovL
AnovL [Deng et al.(2023)Deng, Zhang, Bao, and Li] is, to our knowledge, the only existing test-time adaptation method tailored for zero-shot anomaly localization. It introduces a lightweight adapter on the visual token stream and refines pixel-level anomaly maps by minimizing a self-supervised reconstruction loss; the adapter learns to reconstruct features of normal patches, so that anomalous regions, whose reconstructions incur higher error, become more salient.
To investigate the complementarity of these adaptation schemes, we conducted an ablation in which PILOT’s image-level TTA is combined with AnovL’s pixel reconstruction loss (row “+ AnoVL Loss”). Although this hybrid approach retains PILOT’s adaptive prompt weighting mechanism and the visual adapter architecture, it failed to improve either localization or detection. This result suggests that the reconstruction and pseudo-label losses impose conflicting objectives on the same prompt parameters, generating contradictory update signals that destabilize adaptation and ultimately degrade performance.
By design, AnovL’s adaptation objective targets only pixel-level refinement and does not update any image-level scoring components. As a result, AnovL achieves modest gains in localization accuracy but has no impact on image-level detection (Table 16). In contrast, PILOT’s TTA strategy optimizes a high-confidence, image-level objective by adaptively updating the learnable prompt parameters according to their relevance for each test image, which leads to consistent improvements on both image- and pixel-level metrics. This dual benefit arises because PILOT emphasizes prompt parameters most aligned with the test data, while the fixed attribute embeddings act as semantic anchors to stabilize adaptation and prevent parameter drift.
To investigate the complementarity of these adaptation schemes, we conducted an ablation in which PILOT’s image-level TTA is combined with AnovL’s pixel reconstruction loss (row “+ AnoVL Loss”). Although this hybrid approach retains PILOT’s adaptive prompt weighting mechanism and the visual adapter architecture, it failed to improve either localization or detection. We hypothesize that the reconstruction and pseudo-label losses impose conflicting objectives on the same prompt parameters, driving contradictory update signals that destabilize adaptation and degrade performance.
Table 16: Comparison between PILOT and AnovL on two benchmarks (MVTec AD and VisA). “w/o TTA” and “w/ TTA” indicate performance without and with test-time adaptation, respectively. Arrows denote change under TTA: $\uparrow$(improvement), $\downarrow$(decline), $\circ$(no significant change, $\pm 0.5$).
[…]
Appendix J Details of Attribute Memory Bank Construction
To construct the attribute memory bank, we first define four fixed sets of text templates, each capturing domain-relevant semantics for anomaly detection. Specifically, two sets encode object states: a normal state set, which contains descriptors such as “normal {}” and “flawless {},” and an anomalous state set, which includes terms like “damaged {}” and “imperfect {}.” In addition, we curate three context template lists to capture diverse imaging scenarios: a set for general visual conditions (e.g., “a cropped photo of the {}” or “a blurry photo of a {} for anomaly detection”), an industrial context set (e.g., “an industrial photo of a {}” or “a bright industrial photo of the {}”), and a medical context set (e.g., “a CT scan of a {}” or “an ultrasound image of a {}”).
For each object class, we instantiate all possible combinations by inserting the class name into each entry of the normal and anomalous state sets, and further merging them with every template from the general, industrial, and medical context lists. This combinatorial process results in a comprehensive set of several hundred prompts per class, spanning a wide spectrum of visual conditions and attribute states. For example, for the class “bearing,” we synthesize prompts such as “normal bearing in a cropped photo,” “imperfect bearing in a bright industrial photo,” and “damaged bearing in an ultrasound image.”
All generated prompts are tokenized and encoded using the CLIP text encoder, after which the resulting embeddings are L2-normalized. To facilitate efficient retrieval, the embeddings are partitioned into two disjoint subsets corresponding to normal and anomalous states, thereby enabling the attribute memory bank to provide appropriate semantic vectors during prompt weighting for both image-level and pixel-level anomaly detection tasks. This construction ensures that the memory bank captures a rich and diverse set of attribute semantics, supporting robust prompt weighting and improved generalization to unseen domains.
Appendix K Pseudo Code
Algorithm 1 Training Procedure for PILOT
0: Dataset $D_{aux}=\{(I_{i},y_{i},G_{i})\}$; Learnable Prompt Pool $\mathcal{P}$; Attribute Memory Bank $\mathcal{U}$; Vision Encoder $f_{visual}$; Text Encoder $f_{text}$; epochs $E$; batch size $B$
1:Freeze all weights except: last visual projection, learnable prompt pool, attribute memory bank $\triangleright$ initialize model
2:for $e=1$ to $E$ do
3:  for each batch $(I,y,G)$ in $D_{aux}$ do
4:   $x_{\rm cls},\{X^{(m)}\}_{m\in M}\leftarrow f_{visual}(I)$ $\triangleright$ extract global & patch features from the vision encoder
5:   $\alpha\leftarrow Q_{lrn}(x_{\rm cls},\mathcal{P})$
6:   $\bar{p}^{n},\bar{p}^{a}\leftarrow\sum_{k=1}^{K}\alpha_{k}p_{k}^{n},\sum_{k=1}^{K}\alpha_{k}p_{k}^{a}$ $\triangleright$ aggregate prompts with learned weights
7:   $T_{n},T_{a}\leftarrow f_{text}(\bar{p}^{n},\bar{p}^{a})$ $\triangleright$ encode normal/anomaly prompt
8:
9:   $\beta\leftarrow Q_{att}(x_{\rm cls},\mathcal{U})$
10:   initial Gumbel noise
11:   $\phi_{c}\leftarrow\operatorname{entmax}_{1.5}\bigl{(}\beta_{c}+\text{noise}\bigr{)}$ $\triangleright$ ensure sparsity
12:   $T_{att}\leftarrow\sum_{c=1}^{C}\phi_{c}\,u_{c}$ $\triangleright$ construct attribute representation
13:
14:   $T_{n}^{\rm fus},T_{a}^{fus}\leftarrow\mathrm{Projector}(T_{att},T_{n},T_{a})$ $\triangleright$ fuse by projector
15:   $\hat{y}=\sigma(sim(x_{\rm cls},T_{(\cdot)}^{fus}))$ $\triangleright$ image-level score
16:   $\mathcal{L}_{\rm image}=\mathrm{BCE}(\hat{y},y)$ $\triangleright$ image-level loss
17:   for each $m\in M$ do
18:    $\hat{G}^{(m)}=\sigma(sim(X^{(m)},T_{(\cdot)}^{\rm fus}))$ $\triangleright$ pixel anomaly map
19:   end for
20:   $\mathcal{L}_{\rm pixel}=\frac{1}{|M|}\sum_{m\in M}[\mathrm{Focal}(\hat{G}^{(m)},G)+\mathrm{Dice}(\hat{G}^{(m)},G)]$ $\triangleright$ pixel-level loss
21:   $\mathcal{L}_{\rm PILOT}=\mathcal{L}_{\rm image}+\mathcal{L}_{\rm pixel}$ $\triangleright$ total loss
22:   Update prompt pool, attribute bank, projection using $\mathcal{L}_{\rm PILOT}$
23:  end for
24:end for
In Appendix K, we present high-level pseudocode for both the training and test-time adaptation procedures of PILOT. Algorithm 1 details the end-to-end optimization on the auxiliary dataset $D_{\mathrm{aux}}$, where we jointly update the learnable prompt pool $\mathcal{P}$, attribute memory bank $\mathcal{U}$, and final visual projection layer using combined image-level and pixel-level objectives. Algorithm 2 then describes our confidence-based, image-level objective TTA scheme on an unlabeled target set $D_{\mathrm{tar}}$, in which only the relevant prompt parameters and their associated $r_{k}$, $s_{k}$ embeddings are fine-tuned to align with high-confidence pseudo-labels.
Algorithm 2 Test-Time Adaptation Procedure for PILOT
0: Target dataset $D_{tar}=\{I_{t}\}$; Trained model ($f_{visual}$, $f_{text}$, prompt pool $\mathcal{P}$, attribute bank $\mathcal{U}$); confidence fraction $\rho$; epochs $E$
1:Freeze all weights except selected prompt(s) and associated selector/prototype embeddings
2:for $e=1$ to $E$ do
3:  for each $I_{t}$ in $D_{tar}$ do
4:   $x_{\rm cls}^{t}\leftarrow f_{visual}(I)$ $\triangleright$ extract global features
5:   $\bar{p}^{n},\bar{p}^{a}\leftarrow\sum_{k=1}^{K}\alpha_{k}p_{k}^{n},\sum_{k=1}^{K}\alpha_{k}p_{k}^{a}$ $\triangleright$ aggregate prompts with learned weights
6:   $T_{n},T_{a}\leftarrow f_{text}(\bar{p}^{n},\bar{p}^{a})$ $\triangleright$ encode normal/anomaly prompt
7:   $\beta\leftarrow Q_{att}(x_{\rm cls},\mathcal{U})$
8:   initial Gumbel noise
9:   $\phi_{c}\leftarrow\operatorname{entmax}_{1.5}\bigl{(}\beta_{c}+\text{noise}\bigr{)}$ $\triangleright$ ensure sparsity
10:   $T_{att}\leftarrow\sum_{c=1}^{C}\phi_{c}\,u_{c}$ $\triangleright$ construct attribute representation
11:   $T_{n}^{\rm fus},T_{a}^{fus}\leftarrow\mathrm{Projector}(T_{att},T_{n},T_{a})$ $\triangleright$ fuse by projector
12:   $\tilde{y}=\sigma(\mathrm{sim}(x_{\rm cls}^{t},T_{(\cdot)}^{fus}))$ $\triangleright$ image-level anomaly score
13:  end for
14:  Collect all scores $\{\tilde{y}_{i}\}_{i=1}^{|D_{tar}|}$
15:  Identify top/bottom $\rho$-fraction: $A=\mathrm{Top}_{k}\{\tilde{y}_{i}\}$, $N=\mathrm{Bot}_{k}\{\tilde{y}_{i}\}$, $k=\lceil\rho|D_{tar}|\rceil$
16:  Assign pseudo-labels: $\tilde{y}_{i}=1$ if $i\in A$, $\tilde{y}_{i}=0$ if $i\in N$
17:  $\mathcal{L}_{image}^{TTA}=\frac{1}{|A|+|N|}\sum_{i\in A\cup N}\mathrm{BCE}(\tilde{y}_{i},\text{pseudo-labels})$ $\triangleright$ image-level TTA loss
18:  Update only the prompt(s) and associated embedding relevant to the current input to minimize $\mathcal{L}_{\text{image}}^{\text{TTA}}$
19:end for
Appendix L Visualization
In this section, we present qualitative examples for a subset of categories from each industrial dataset generated by PILOT. Figures 4-25 illustrate representative images and corresponding anomaly maps; to maintain clarity, we omit exhaustive visualizations of every class.
Figure 4: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the bottle category in the MVTec AD dataset. All results are generated by PILOT.
Figure 5: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the grid category in the MVTec AD dataset. All results are generated by PILOT.
Figure 6: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the metal nut category in the MVTec AD dataset. All results are generated by PILOT.
Figure 7: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the wood category in the MVTec AD dataset. All results are generated by PILOT.
Figure 8: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the cable category in the MVTec AD dataset. All results are generated by PILOT.
Figure 9: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the chewinggum category in the VisA dataset. All results are generated by PILOT.
Figure 10: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the cashew category in the VisA dataset. All results are generated by PILOT.
Figure 11: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the pcb4 category in the VisA dataset. All results are generated by PILOT.
Figure 12: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the pcb2 category in the VisA dataset. All results are generated by PILOT.
Figure 13: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the 01 category in the BTAD dataset. All results are generated by PILOT.
Figure 14: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the 02 category in the BTAD dataset. All results are generated by PILOT.
Figure 15: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the 03 category in the BTAD dataset. All results are generated by PILOT.
Figure 16: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the electrical commutators category in the SDD dataset. All results are generated by PILOT.
Figure 17: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the class 1 category in the DAGM dataset. All results are generated by PILOT.
Figure 18: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the class 2 category in the DAGM dataset. All results are generated by PILOT.
Figure 19: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the class 3 category in the DAGM dataset. All results are generated by PILOT.
Figure 20: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the class 4 category in the DAGM dataset. All results are generated by PILOT.
Figure 21: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the Blotchy category in the DTD dataset. All results are generated by PILOT.
Figure 22: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the Fibrous category in the DTD dataset. All results are generated by PILOT.
Figure 23: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the Woven category in the DTD dataset. All results are generated by PILOT.
Figure 24: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the Matted category in the DTD dataset. All results are generated by PILOT.
Figure 25: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the Perforated category in the DTD dataset. All results are generated by PILOT.